OpenStudy (anonymous):

y varies jointly with p and q. When p = 3 and q = 3, y = 90. Find y when p = 7 and q = 5

OpenStudy (anonymous):

i forgot the whole y equals k something whatever formula....

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

y = kpq

OpenStudy (anonymous):

so k=10 and y=350?

OpenStudy (anonymous):

y/(pq) = 90/(3)(3) = y/(7)(5) Solve for y

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

k = 10, but y NOT = 350

OpenStudy (anonymous):

yes it is?

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

yes, you are right.

OpenStudy (anonymous):

why u isa confussin meh -.-

OpenStudy (anonymous):

Don't understand. Speak English.

OpenStudy (anonymous):

das is engrish

OpenStudy (anonymous):

isa i use da difrent dialect

OpenStudy (anonymous):

Have a good day. Bye now.

OpenStudy (anonymous):

lol....poor kid..so confused....im sorry im kinda hyper lol